Hello,
I have a question I need help with. Your time is appreciated.
A college student earned $5500 during summer vacation working as a waiter in a popular restaurant. The student invested part of the money at 9% and the rest at 7%. If the student received a total of $429 in interest at the end of the year, how much was invested at 9%?
Thank you

Answer by stanbon(75887)

A college student earned $5500 during summer vacation working as a waiter in a popular restaurant.
The student invested part of the money at 9% and the rest at 7%.
If the student received a total of $429 in interest at the end of the year, how much was invested at 9%?
---
Using 2 variables you get 2 equations:
n + s = 5500 dollars
---
0.09n + 0.07s = 429 dollars
-----
Multiply thru the 1st equation by 7
Multiply thru the 2nd equation by 100
-----
7n + 7s = 7*5500
9n + 7s = 42900
-------------------
Subtract and solve for "n":
2n = 4400
n = $2200 (amt invested at 9%)
==================================
Cheers,
Stan H.
============